How do I solve this equation below by using the square root property?
3(x+5)^2=54
The solution set is {}?
(type an exact answer, using radicals as needed.

Answer by stanbon(75887)

solve this equation below by using the square root property?
3(x+5)^2=54
(x+5)^2 = 18
-------
x+5 = sqrt(18) or x+5 = -sqrt(18)
-----
x = -5+3sqrt(2) or x = -5-3sqrt(2)
====================================
Cheers,
Stan H.
==============